poppie Bonds have a 6.26% coupon rate and pay interest semiannually currently the bond are quoted at 114.4 immature in four years what is the yield to maturity
Answer #1

Given: B=114.4 t = 4 years coupn= 6-26% = 6.26 let yield to maturity - y 114.4= 6-26 + 6.26 & 6.26 & 100+ 6.26 solver: sowing

Please do rate me and mention doubts in the comments section.